OpenStudy (anonymous):

Help Please<3 Activity 1. 1. a. What is the common ratio of the sequence 2, 4, 8, 16,...? b. What are the next three terms in the sequence? 2. a. What is the common ratio of the sequence 80, 20, 5, 5/4, ...? b. What are the next three terms in the sequence? 3. a. What is the common ratio of the sequence 2, -6, 18, -54, ...? b. What are the next three terms in the sequence? Thank You <3

OpenStudy (haseeb96):

1.r=4/2 = 2 is the common ratio So the next three terms will be by multiply 2 with the last term

OpenStudy (haseeb96):

16×2=32 32×2=64 64×2=128 These are next three terms

OpenStudy (anonymous):

Thank You! @Haseeb96

OpenStudy (anonymous):

What about 2 & 3? @Haseeb96

OpenStudy (haseeb96):

R=-6/2 =-3 so the common ratio is -3 when u multiply -3 with the last term u will get the next term and then again multiply -3 with the last up conning term and get the 2nd term and again mulitly like and get 3rd term

OpenStudy (haseeb96):

-54x-3=? 162×-3=? Same for third term